OpenStudy (princess123):

The area of the rectangular playground enclosure at South school is 500 square meters. The length of the playground is 5 meters longer than the width. Finf the dimensions of the playground, in width. width=? meters length=? meters

OpenStudy (anonymous):

l*w = 500 l = 5+w w^2 +5w = 500 (w+25)(w-20) = 0 w must be greater than 0, so w = 20, l = 25

OpenStudy (princess123):

thanks!